Exercise 3-6 Preparing adjusting entries LO P1 a. Depreciation on the companys equipment for 2017 is computed to be $11,000.a. Depreciation on the companys equipment for 2017 is computed to be $11,000. b. The Prepaid Insurance account had a $7,000a. Depreciation on the companys equipment for 2017 is computed to be $11,000. b. The Prepaid Insurance account had a $7,000a. Depreciation on the companys equipment for 2017 is computed to be $11,000. b. The Prepaid Insurance account had a $7,000a. Depreciation on the companys equipment for 2017 is computed to be $11,000. b. The Prepaid Insurance account had a $7,000a. Depreciation on the companys equipment for 2017 is computed to be $11,000 b. The Prepaid Insurance account had a $7,000 d
Answer #1
General Journal Debit Credit
a Depreciation expense—Equipment 11000
       Accumulated depreciation—Equipment 11000
b Insurance expense 5490 =7000-1510
       Prepaid insurance 5490
c Office supplies expense 2639 =230+2680-271
       Office supplies 2639
d Unearned fee revenue 5000 =15000*1/3
     Fee revenue 5000
e Insurance expense 3690
      Prepaid insurance 3690
f Wages expense 6000
       Wages payable 6000